OpenStudy (anonymous):

Which polynomial is a difference of two squares? x2 + 25 x2 − 25 x2 − 10 x2 + 10 i dont have idea i need help ...

OpenStudy (tennistar):

2

OpenStudy (anonymous):

x^2 - 25

OpenStudy (tennistar):

both are squares and are being subtracted

OpenStudy (anonymous):

hence the word difference ^

OpenStudy (jackellyn):

It has to be in the form a^2 - b^2 a=1^2 =1 b=5^2 =5 x^2-25

OpenStudy (anonymous):

oh i think i get it =D